Question 19 Let X denote the number of bits received in error in a digital communication channel, and assume that X is a binomial with p = 0.0003. If 1000 bits are transmitted, determine the following Round your answers to four decimal places (e.g. 98.7654) (a) P(X = 1) (b) P(X 2 1) (c) P(X s 2) (d) mean and variance of X. (d) Mean = variance =
Answer #1

Given X:- Mumber of bits recorded in enor nr-1000 I p= 0.0003 XN Binomial Plx - 1) (n=1000, P=0.0003) a) (1000) (0.00132 (-0.c) PCX of 22) = P(x=0) + P(x=1) + P(x-2) - 0.7408 4 0.2223 + love (0.6003) (1-0-000396 - 2 – 0.0333 0.7408 +0.2223+ 0.9964 =